Grade 11Mechanics

Two identical point masses, eah ofmass M are connected to one another bya massless string of length L. A constant force F is applied at the mid point of the string. If l be the instantneous distance between the two masses. What will be the acceleration of each mass.

SAI SWARAJ SHAW
9 Years agoGrade 11

1 Answer

Rituraj Tiwari
5 Years ago

To analyze the scenario with two identical point masses connected by a massless string and subjected to a constant force, we need to break down the forces acting on each mass and apply Newton's second law of motion. This will help us determine the acceleration of each mass given that they are connected and influenced by the same force.

Basic Setup and Forces Involved

We have two masses, each with a mass M, connected by a string of length L. When a constant force F is applied at the midpoint of the string, the force effectively acts on both masses because they are linked through the string.

Identifying the Total Force

Since the string is massless, we can assume that the tension within the string is uniform. The important point here is that both masses will experience the same tension in the string, which will affect their acceleration. The total force applied F is split evenly between the two masses due to their identical nature and the fact that they are symmetrically positioned.

Applying Newton's Second Law

Newton's second law states that the net force acting on an object is equal to the mass of the object multiplied by its acceleration (F = ma). In our case, we can analyze one of the masses:

  • For mass 1: The net force acting on mass 1 is the tension T pulling it towards mass 2, and the net force is also equal to the force F applied to the system divided by 2 (since both masses share the force equally).
  • For mass 2: A similar analysis applies, where it experiences tension T pulling it towards mass 1.

Calculating the Acceleration

Let’s denote the acceleration of each mass as 'a'. For each mass, we can write the equation as follows:

For mass 1: T = M * a

For mass 2: T = M * a

Since the total force F is responsible for accelerating the entire system of two masses, we can express this as:

F = 2 * T

Replacing T with Ma in the equation, we have:

F = 2 * (M * a)

This simplifies to:

a = F / (2M)

Conclusion on Acceleration

Therefore, the acceleration of each mass is determined by the equation a = F / (2M). This means that the acceleration of each mass is directly proportional to the applied force and inversely proportional to twice their individual mass. This outcome is crucial for understanding how forces distribute in a system involving connected objects.

Real-World Analogy

Imagine two people holding a rope with a weight in the middle. If one person pulls on the rope, both will feel the tension and will accelerate together. The same principle applies here; the force F acts on the system, causing both masses to accelerate uniformly.